OpenStudy (anonymous):

What are the real numbers of x and y that makes the equation true?(i = root -1) 2 + 3i = x + yi Can you explain the answer? I already know the answer, but I don't know how you can get it.

OpenStudy (anonymous):

isnt the answer x=2 and y =3 ? just to make sure

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

well when i look at this it seems pretty straight forward like: a + bi = x + yi a = x bi = yi => b =y

OpenStudy (anonymous):

i is treated somewhat like a variable x but when you get i^2 it becomes -1

OpenStudy (anonymous):

ohhh thank you :)